# CECILIA: Gas-Phase Physical Conditions and Multi-Element Chemistry at Cosmic Noon

### v1

@@ 4.1 Total O Abundance
Figure 6 also reveals the presence of relatively metal-rich galaxies in the CECILIA sample. The left panel plots the M. Asplund et al. (2021) solar O/H ratio and its uncertainty as a green dashed line and shaded area, respectively. Of the 19 galaxies plotted, 5 are consistent with the solar O/H666This excludes RK120, a galaxy with direct 12+log(O/H) 9.0 and other peculiar abundance patterns. The MOSFIRE cross-band slit-loss corrections for RK120 are poorly constrained, which affects the inferred reddening and strong-line ratios that get propagated into the Te and abundances. We do not discuss RK120 in the remainder of our analysis.. Such high metallicities are uncommon in high-z galaxies. For example, the highest gas-phase metallicity measured in the EXCELS SFGs is 12+log(O/H)8.380.30 at z4.234 (T. M. Stanton et al., 2025), which is consistent with the upper metallicity measured in the CEERS (S. L. Finkelstein et al., 2025) auroral line sample in R. L. Sanders et al. (2024a). Despite the growing sample of high-z galaxies with direct Te and abundances, including 16 and 41 SFGs at z1.3 reported in E. Cataldi et al. (2025) and R. L. Sanders et al. (2025), respectively, there are only three galaxies in the literature with solar O/H abundance in the ISM. The number is more than doubled with CECILIA, and the two CECILIA galaxies with 12+log(O/H)8.70 dex constitute the largest direct-method metallicities measured with JWST to date.
@@
In general, it is challenging to apply the direct abundance method at solar O/H, as metal CEL emission is the dominant cooling mechanism in the ISM. As metallicity increases, collisional excitation from free electrons and spontaneous emission from the abundant metals acts to remove energy from the gas and significantly lower the overall temperature of the ISM. At these low Te, emission from metal fine-structure lines in the infrared becomes more significant as the intensity of the optical CELs decreases (M. A. Dopita & R. S. Sutherland, 2003). The decrease in intensity is especially notable for the optical Te-sensitive auroral lines, which originate from energy levels with high excitation energies ranging from 3.36 eV for [S III]6314 to 5.35 eV for [O III]4364. Even from large samples of metal-rich nebulae in the local Universe, there are relatively few H II regions with direct-method 12+log(O/H)8.69 dex (see compilation and discussion in J. E. Méndez-Delgado et al., 2024). Given these trends, it is notable that roughly a quarter of the CECILIA galaxies with direct O/H, including two with multiple Te measurements, show O/H consistent with the solar abundance ratio.
@@
To assess if the direct O/H measurements are consistent with the physical conditions and emission line properties of the CECILIA galaxies, we first consider the Te distributions plotted in Figure 5. The median Te[O II] measured in the CECILIA galaxies is 10700 K, similar to the Te[O II] measured in the metal-rich CHAOS H II regions and less than the Te[O II] measured in the metal-poor EELGs at z0. We then compare the O/H abundances from the CECILIA and CHAOS samples to determine if CECILIA is offset to higher metallicities than anticipated from the Te distribution. Considering the CHAOS H II regions with Te[O II] 10700500 K, the average 12+log(O/H) in the 20 regions is 8.38 dex with a standard deviation of 0.15 dex. This is comparable with the average 12+log(O/H) in the CECILIA galaxies, 8.28 dex, indicating that the combination of optical CELs and direct Te produces O/H that is consistent with nebulae of similar Te conditions.
@@
Next, we reproduce the N2-BPT in the right panel of Figure 6, where we have focused on the Te sample and color-coded each point by the direct O/H. The observed trend matches the expected behavior of galaxies in this line ratio diagram, which is sensitive to ionization, total O/H, and the relative abundance of N/O. At low-O/H, the highly-ionized gas cools via [O III]5008 emission and results in large [O III]/H, while low N abundance and weak low-ionization emission lines produce in low [N II]/H. As metallicity increases, so too do the available coolants in the form of O and secondary N, while the hardness of the ionizing spectrum decreases. This combination leads to lower [O III]/H as [N II]/H increases. In CECILIA, the galaxies with log([N II]/H) 1.4 and log([N II]/H) 1.0 have average 12+log(O/H) of 8.04 and 8.48 dex, respectively, in qualitative agreement with the predicted evolution in the BPT diagram.
@@
Systematic uncertainties could bias the reported O abundances in CECILIA, the most plausible being the use of the [O II] nebular lines from the Keck/MOSFIRE spectrum to derive Te[O II] (see discussion in §2.6). One method to assess this is to use the Te[S III] measurement and a Te scaling relation to infer the low-ionization zone temperature, then apply that temperature with the [O II] auroral lines to measure O+ from the NIRSpec data alone. This method was adopted for the chemical abundance study of D40 (N. S. J. Rogers et al., 2024), and is only possible for the galaxies with simultaneous [S III]6314 and [O II]7322,32 detections (i.e., those plotted in Figure 5). Of the four galaxies (i.e., excluding two with direct Te[N II] where Te[O II] is not adopted), the use of Te[S III] alone results in statistically consistent O/H. While we acknowledge that some galaxies could be affected by a bias in Te[O II], this exercise also relies on the shape of the Te[O II]-Te[S III] scaling relation that is, presently, not well constrained on empirical Te data at high-z. The Te scaling relations are another source of systematic uncertainty, particularly for the high-ionization zone Te that must be inferred for the majority of CECILIA galaxies. Nonetheless, these consistency checks indicate that the O/H in the CECILIA galaxies is not uncharacteristic of galaxies with similar Te or CEL emission, suggesting that SFGs at Cosmic Noon could reach solar O/H in the gas phase.
@@
We now compare the Te and O/H measurements in the CECILIA galaxies to other high-z SFGs with published direct-method Te and metallicities acquired with JWST. In Figure 7, we plot 12+log(O/H) against Te[O II] (or Te[N II], when available), Te[S III], and Te[O III] for the CECILIA sample and 85 literature SFGs at z1.3 (from B. Welch et al., 2024, 2025; T. Morishita et al., 2025; T. M. Stanton et al., 2025; K. Z. Arellano-Córdova et al., 2025; R. L. Sanders et al., 2024a, 2025; E. Cataldi et al., 2025). In each panel, we denote whether each Te is directly measured or inferred from a Te scaling relation. For CECILIA, the direct Te are plotted as yellow circles (or yellow squares for Te[N II]) while the inferred Te are represented by red circles. The purple and white triangles indicate the literature SFGs with direct and inferred Te, respectively. When an inferred Te[O II] or Te[S III] is not reported, we apply the D. R. Garnett (1992) scaling relations along with the reported Te[O III] to include the galaxy in Figure 7. Since there are so few Te[O II] and Te[S III] samples at z1, most of the literature Te in the left and middle panels are inferred from the reported Te[O III].
@@
As expected, there is a strong anti-correlation between gas-phase metallicity and Te, although Figure 7 provides additional insight into the state of direct Te and abundances in the era of JWST. Te[O III] is the most common temperature measured at high-z and ranges from 0.85-2.5104 K. The resulting metallicities, 7.012+log(O/H)8.5, are limited by the inherent challenge of applying the direct abundance method with Te[O III] alone. When the ISM is highly ionized and at low-metallicity, O2+ is the dominant ionization state of O and the ISM is sufficiently hot to permit auroral [O III] excitation. However, in the same environments the [O II] and [S III] auroral lines are relatively faint owing to the scarcity of O+ and S in the ISM. This physical scenario manifests as the dearth of direct Te[O II] at 12+log(O/H)8.0 in the literature sample, and the lack of literature Te[S III] likely originates from a combination of exceptionally faint auroral emission in tandem with insufficient wavelength coverage for the nebular [S III] lines.
@@
In contrast, low- and intermediate-ionization state auroral lines are easier to detect in metal-rich or low-ionization environments where [O III]4364 becomes prohibitively faint (e.g., N. S. J. Rogers et al., 2021). Accordingly, most direct Te[O II] in the literature SFGs occur between 12+log(O/H) of 8.0 and 8.5. Measuring Te[O II] in the cooler, metal-rich ISM becomes challenging owing to the diminishing intensity of the [O II] auroral lines with decreasing temperature. Detecting the relatively faint [O II] and [S III] auroral lines over a wider range of O/H is possible from the ultra-deep CECILIA observations, and these data extend the O/H-Te anti-correlation to higher metallicities than the literature SFGs.
@@ 4.2 S/O and Ar/O
In local nebulae, the abundance ratios of S/O and Ar/O in the ISM tend to be constant and very close to the solar abundance ratios (Y. I. Izotov et al., 2006; D. A. Berg et al., 2020; A. Amayo et al., 2021; N. S. J. Rogers et al., 2022; K. Z. Arellano-Córdova et al., 2024; C. Esteban et al., 2025). This trend suggests a similar production mechanism for all three elements, that being the process in the interiors of massive stars. However, the Type Ia SNe yield models of C. Kobayashi et al. (2020) indicate that S and Ar can be produced in excess of O in Type Ia SNe. If the abundance of S and Ar are sensitive to enrichment from Type Ia SNe, then the ratios of S/O and Ar/O should evolve based on the relative enrichment of these two sources and the SFH of a galaxy. Variations in the Ar/O abundance have been observed in local nebulae, such as the planetary nebulae of M31 (M. Arnaboldi et al., 2022; C. Kobayashi et al., 2023). Recently, S. Bhattacharya et al. (2025b) reanalyzed SFGs in SDSS and uncovered a subtle evolution in the O/Ar ratio with Ar/H and stellar mass, where high-mass galaxies show larger O/Ar at fixed Ar/H relative to low-mass galaxies. This trend reflects the lower star-formation efficiency (SFE) and gradual Type Ia enrichment in low-mass galaxies that contributes to the decrease of O with respect to Ar or other Type Ia nucleosynthetic products.
@@
The chemical abundance patterns in galaxies at Cosmic Noon, which are undergoing recent and rapid star formation, may be sensitive to the relative enrichment from CCSNe and Type Ia SNe. In Figure 8, we plot the direct S/O and Ar/O abundance ratios measured from the full CECILIA sample against the corresponding gas-phase O/H measured in each galaxy. For comparison, the S/O and Ar/O abundance patterns in the CHAOS H II regions and low-metallicity EELGs at z0 are plotted as gray diamonds and blue squares, respectively. We also provide the solar abundance ratios as the circle dot and dashed green line. The CECILIA galaxies plotted in Figure 8 represent the largest sample of S and Ar abundances in galaxies at z2 compiled from JWST observations, which is enabled because of the design of CECILIA. Reliable S abundances require the detection of the [S III] nebular lines in the rest-NIR (see discussion in A. Amayo et al., 2021), while the [Ar III]7138 emission line is faint (2% the flux of H in the CECILIA galaxies) owing to the scarcity of Ar in the ISM. As such, deep spectroscopy required to measure [Ar III]7138 and the [S III] auroral line must be matched with broad wavelength coverage to detect the nebular [S III] lines and derive a reliable Te[S III].
@@
These abundance trends reveal that the typical galaxy in the CECILIA sample has lower S/O and Ar/O than the solar abundance ratio. We measure an average log(S/O) 1.770.21 dex and log(Ar/O) 2.630.12 dex (plotted as dot dashed red lines in Figure 8), which are 0.20 and 0.32 dex lower than the respective solar ratios. Consistent with local chemical abundance surveys, there is no apparent evolution in S/O or Ar/O with total metallicity. From CECILIA, there is growing evidence for non-solar abundance patterns in galaxies during the epoch of Cosmic Noon. Given the contribution of Type Ia SNe to the gas-phase S and Ar abundances, a natural explanation for the enhanced O relative to these elements is the preferential enrichment from CCSNe. We provide the S/O and Ar/O abundance ratios predicted from the IMF-averaged CCSNe yields for metal-poor stars from C. Kobayashi et al. (2006) as black dotted lines in Figure 8. We observe that the average S/O and Ar/O measured in CECILIA is in good agreement with the yield predictions, further supporting a scenario of significant CCSNe enrichment.
@@
The average ratios measured in the full sample are in good agreement with the abundance trends in D40, the first galaxy with direct Te[S III], S and Ar abundance measurements at z2 (N. S. J. Rogers et al., 2024). As discussed in that work, certain systematic uncertainties could produce an offset to low S/O and Ar/O. First, D40 was previously missing a direct measurement of the low-ionization zone Te, such that the earlier O+ and S+ abundances could have been biased. From the combined MOSFIRE and NIRSpec data reported here, the Te[O II] in D40 agrees with the Te inferred from the other direct temperatures. While the lack of Te information did not appear to significantly bias the abundance ratios in D40, galaxies with a single Te are reliant on the functional form of the adopted Te scaling relations. Second, the selected ICFs could fail to account for all unobserved ionization states of S and Ar in the ISM. However, the trend of sub-solar S/O and Ar/O persists when the analysis is repeated with the A. Amayo et al. (2021) ICFs, indicating that a substantial change to the shape of the ICFs is required to bring the ratios into agreement with the solar abundances (see also T. M. Stanton et al., 2025). Finally, the use of the [O II] nebular lines from the MOSFIRE spectra could bias to larger O+ abundance (see discussion in previous subsection). However, when we repeat the analysis with only Te[S III] and use the [O II] auroral lines for the O+ abundance (similar to the original analysis for D40), we find that the S/O abundances in five of the six galaxies remain consistent within uncertainties. For BX628, which shows a larger S/O than the solar ratio in Figure 8, using Te[S III] to infer the low-ionization zone Te produces a significantly lower S/O that is in good agreement with the CECILIA average. While we cannot rule out the potentially distinct S or Ar ICF, the full CECILIA sample spans a range in ionization such that the bulk S/O and Ar/O trends should be representative of the chemical abundance patterns in typical SFGs at Cosmic Noon.
@@
Currently, there are relatively few high-z galaxies with direct S/O abundances owing to the observational challenges associated with detecting the NIR [S III] lines. To compare with existing abundance measurements in Figure 9, we focus on Ar/O measured in other high-z galaxy surveys. We include Ar/O from the observations of SGAS 1723+34 from the TEMPLATES survey (B. Welch et al., 2024), the Sunburst Arc (B. Welch et al., 2025), MACSJ1149-WR1 (T. Morishita et al., 2025), the EXCELS survey from T. M. Stanton et al. (2025), and the compilation of Ar/O from S. Bhattacharya et al. (2025a) excluding D40 and SGAS 1723+34. We use the published Ar/O in each source except for MACSJ1149-WR1, which is reported to have log(Ar/O) 2.400.02 dex. The reported uncertainty on Ar/O is likely underestimated, considering that the uncertainty in the total O abundance is 0.1 dex and the low-ionization zone temperature has uncertainty in excess of 3000 K. We recompute the Ar/O abundance in MACSJ1149-WR1 using the reported line intensities and ionization zone temperatures in T. Morishita et al. (2025) and find log(Ar/O) 2.390.22 dex, which we use when comparing to the CECILIA sample.
@@
Figure 9 shows log(Ar/O) vs. redshift for 36 galaxies at 1.2z5.8, where the literature abundances are plotted as purple triangles. The Ar/O abundances in the literature galaxies are generally at or below the solar abundance ratio. To assess the average trends of these high-z galaxies, we measure the average Ar/O ratio in three redshift bins: 1.5z2.5, 2.5z3.5, and 3.5z4.5. Each bin has more than five galaxies with direct abundances from which to measure the average Ar/O, which are plotted as white stars in Figure 9. While the highest redshift bin has the largest uncertainty, the galaxies at 1.5z2.5 and 2.5z3.5 show similar log(Ar/O) and are in agreement with the CCSNe yield predictions. Excluding the CECILIA sample, the number of galaxies in the 1.5z2.5 and 2.5z3.5 bins decrease from 15 to 5 and 12 to 4, respectively. Without the CECILIA galaxies, the average log(Ar/O) in these bins increases to log(Ar/O) 2.48 dex, consistent with the binned averages in Figure 9 and over 0.15 dex below the solar Ar/O ratio.

### v2

@@ 4.1 Total O Abundance
Figure 6 also reveals the presence of relatively metal-rich galaxies in the CECILIA sample. The left panel plots the asplund2021 solar O/H ratio and its uncertainty as a green dashed line and shaded area, respectively. Of the 19 galaxies plotted, 6 are consistent with the solar O/H666This excludes RK120, a galaxy with direct 12+log(O/H) 9.0 and other peculiar abundance patterns. The MOSFIRE cross-band slit-loss corrections for RK120 are poorly constrained, which affects the inferred reddening and strong-line ratios that get propagated into the Te and abundances. We do not discuss RK120 in the remainder of our analysis.. Such high metallicities are uncommon in high-z galaxies. For example, the highest gas-phase metallicity measured in the EXCELS SFGs is 12+log(O/H)8.380.30 at z4.234 (stanton2025), which is consistent with the upper metallicity measured in the CEERS (finkelstein2025) auroral line sample in sanders2024sl. Despite the growing sample of high-z galaxies with direct Te and abundances, including 16 and 41 SFGs at z1.3 reported in cataldi2025 and sanders2025, respectively, there are only three galaxies in the literature with solar O/H abundance in the ISM. The number is more than doubled with CECILIA, and the three CECILIA galaxies with 12+log(O/H)8.70 dex constitute the largest direct-method metallicities measured with JWST to date.
@@
In general, it is challenging to apply the direct abundance method at solar O/H, as metal CEL emission is the dominant cooling mechanism in the ISM. As metallicity increases, collisional excitation from free electrons and spontaneous emission from the abundant metals acts to remove energy from the gas and significantly lower the overall temperature of the ISM. At these low Te, emission from metal fine-structure lines in the infrared becomes more significant as the intensity of the optical CELs decreases (dopita2003). The decrease in intensity is especially notable for the optical Te-sensitive auroral lines, which originate from energy levels with high excitation energies ranging from 3.36 eV for [S III]6314 to 5.35 eV for [O III]4364. Even from large samples of metal-rich nebulae in the local Universe, there are relatively few H II regions with direct-method 12+log(O/H)8.69 dex (see compilation and discussion in mendez-delgado2024). Given these trends, it is notable that roughly a quarter of the CECILIA galaxies with direct O/H, including two with multiple Te measurements, show O/H consistent with the solar abundance ratio.
@@
To assess if the direct O/H measurements are consistent with the physical conditions and emission line properties of the CECILIA galaxies, we first consider the Te distributions plotted in Figure 5. The median Te[O II] measured in the CECILIA galaxies is 10700 K, similar to the Te[O II] measured in the metal-rich CHAOS H II regions and less than the Te[O II] measured in the metal-poor EELGs at z0. We then compare the O/H abundances from the CECILIA and CHAOS samples to determine if CECILIA is offset to higher metallicities than anticipated from the Te distribution. Considering the CHAOS H II regions with Te[O II] 10700500 K, the average 12+log(O/H) in the 24 regions is 8.41 dex with a standard deviation of 0.14 dex. This is comparable with the average 12+log(O/H) in the CECILIA galaxies, 8.29 dex, indicating that the combination of optical CELs and direct Te produces O/H that is consistent with nebulae of similar Te conditions.
@@
Next, we reproduce the N2-BPT in the right panel of Figure 6, where we have focused on the Te sample and color-coded each point by the direct O/H. The observed trend matches the expected behavior of galaxies in this line ratio diagram, which is sensitive to ionization, total O/H, and the relative abundance of N/O. At low-O/H, the highly-ionized gas cools via [O III]5008 emission and results in large [O III]/H, while low N abundance and weak low-ionization emission lines result in low [N II]/H. As metallicity increases, so too do the available coolants in the form of O and secondary N, while the hardness of the ionizing spectrum decreases. This combination leads to lower [O III]/H as [N II]/H increases. In CECILIA, the galaxies with log([N II]/H) 1.4 and log([N II]/H) 1.0 have average 12+log(O/H) of 8.02 and 8.48 dex, respectively, in qualitative agreement with the predicted evolution in the BPT diagram. Finally, an evolution in O/H is expected from the mass-metallicity relation (MZR, lequeux1979; tremonti2004). The average stellar masses of the CECILIA galaxies with 12+log(O/H) 8.5 dex and 7.5 dex are log(M∗/M⊙) 9.9 and 9.3, respectively. This is in qualitative agreement with the trends of other high-z galaxies (e.g., erb2006metal; sanders2021; strom2022; curti2024), although we defer a detailed study of the MZR at Cosmic Noon for a future work (see raptis2025b).
@@
Systematic uncertainties could bias the reported O abundances in CECILIA, the most plausible being the use of the [O II] nebular lines from the Keck/MOSFIRE spectrum to derive Te[O II] (see discussion in §2.6). One method to assess this is to use the Te[S III] measurement and a Te scaling relation to infer the low-ionization zone temperature, then apply that temperature with the [O II] auroral lines to measure O+ from the NIRSpec data alone. This method was adopted for the chemical abundance study of D40 (rogers2024), and is only possible for the galaxies with simultaneous [S III]6314 and [O II]7322,32 detections (i.e., those plotted in Figure 5). Of the four galaxies (i.e., excluding two with direct Te[N II] where Te[O II] is not adopted), the use of Te[S III] alone produces O/H consistent with the reported values within 1, with an average difference of 0.03 dex. While we acknowledge that some galaxies could be affected by a bias in Te[O II], this exercise also relies on the shape of the Te[O II]-Te[S III] scaling relation that is, presently, not well constrained on empirical Te data at high-z. The Te scaling relations are another source of systematic uncertainty, particularly for the high-ionization zone Te that must be inferred for the majority of CECILIA galaxies. Nonetheless, these consistency checks indicate that the O/H in the CECILIA galaxies is not uncharacteristic of galaxies with similar Te or CEL emission, suggesting that SFGs at Cosmic Noon could reach solar O/H in the gas phase.
@@
We now compare the Te and O/H measurements in the CECILIA galaxies to other high-z SFGs with published direct-method Te and metallicities acquired with JWST. In Figure 7, we plot 12+log(O/H) against Te[O II] (or Te[N II], when available), Te[S III], and Te[O III] for the CECILIA sample and 85 literature SFGs at z1.3 (from welch2024; welch2025; morishita2025; stanton2025; arellano-cordova2025; sanders2024sl; sanders2025; cataldi2025). In each panel, we denote whether each Te is directly measured or inferred from a Te scaling relation. For CECILIA, the direct Te are plotted as yellow circles (or yellow squares for Te[N II]) while the inferred Te are represented by red circles. The purple and white triangles indicate the literature SFGs with direct and inferred Te, respectively. When an inferred Te[O II] or Te[S III] is not reported, we apply the garnett1992 scaling relations along with the reported Te[O III] to include the galaxy in Figure 7. Since there are so few Te[O II] and Te[S III] samples at z1, most of the literature Te in the left and middle panels are inferred from the reported Te[O III].
@@
As expected, there is a strong anti-correlation between gas-phase metallicity and Te (see also flury2020), although Figure 7 provides additional insight into the state of direct Te and abundances in the era of JWST. Te[O III] is the most common temperature measured at high-z and ranges from 0.85-2.5104 K. The resulting metallicities, 7.012+log(O/H)8.5, are limited by the inherent challenge of applying the direct abundance method with Te[O III] alone. When the ISM is highly ionized and at low-metallicity, O2+ is the dominant ionization state of O and the ISM is sufficiently hot to permit auroral [O III] excitation. However, in the same environments the [O II] and [S III] auroral lines are relatively faint owing to the scarcity of O+ and S in the ISM. This physical scenario manifests as the dearth of direct Te[O II] at 12+log(O/H)8.0 in the literature sample, and the lack of literature Te[S III] likely originates from a combination of exceptionally faint auroral emission in tandem with insufficient wavelength coverage for the nebular [S III] lines.
@@
In contrast, low- and intermediate-ionization state auroral lines are easier to detect in metal-rich or low-ionization environments where [O III]4364 becomes prohibitively faint (e.g., rogers2021). Accordingly, most direct Te[O II] in the literature SFGs occur between 12+log(O/H) of 8.0 and 8.5. Measuring Te[O II] in the cooler, metal-rich ISM becomes challenging owing to the diminishing intensity of the [O II] auroral lines with decreasing temperature. Detecting the relatively faint [O II] and [S III] auroral lines over a wider range of O/H is possible from the ultra-deep CECILIA observations, and these data extend the O/H-Te anti-correlation to Te8000 K, consistent with the lowest Te and highest direct-method metallicities measured with JWST.
@@ 4.2 S/O and Ar/O
In local nebulae, the abundance ratios of S/O and Ar/O in the ISM tend to be constant and very close to the solar abundance ratios (izotov2006; berg2020; amayo2021; rogers2022; esteban2025). This trend suggests a similar production mechanism for all three elements, that being the process in the interiors of massive stars. However, the Type Ia SNe yield models of kobayashi2020SNIa indicate that S and Ar can be produced in excess of O in Type Ia SNe. If the abundance of S and Ar are sensitive to enrichment from Type Ia SNe, then the ratios of S/O and Ar/O should evolve based on the relative enrichment of these two sources and the SFH of a galaxy. Variations in the Ar/O abundance have been observed in local nebulae, such as the planetary nebulae of M31 (arnaboldi2022; kobayashi2023). Recently, bhattacharya2025 reanalyzed SFGs in SDSS and uncovered a subtle evolution in the O/Ar ratio with Ar/H and stellar mass, where high-mass galaxies show larger O/Ar at fixed Ar/H relative to low-mass galaxies. This trend reflects the lower star-formation efficiency (SFE) and gradual Type Ia enrichment in low-mass galaxies that contributes to the decrease of O with respect to Ar or other Type Ia nucleosynthetic products.
@@
The chemical abundance patterns in galaxies at Cosmic Noon, which are undergoing recent and rapid star formation, may be sensitive to the relative enrichment from CC and Type Ia SNe. In Figure 8, we plot the direct S/O and Ar/O abundance ratios measured from the full CECILIA sample against the corresponding gas-phase O/H measured in each galaxy. For comparison, the S/O and Ar/O abundance patterns in the CHAOS H II regions and low-metallicity EELGs at z0 are plotted as gray diamonds and blue squares, respectively. We also provide the solar abundance ratios as the circle dot and dashed green line. The CECILIA galaxies plotted in Figure 8 represent the largest sample of S and Ar abundances in galaxies at z2 compiled from JWST observations, which is enabled because of the design of CECILIA. Reliable S abundances require the detection of the [S III] nebular lines in the rest-NIR (see discussion in amayo2021), while the [Ar III]7138 emission line is faint (2% the flux of H in the CECILIA galaxies) owing to the scarcity of Ar in the ISM. As such, deep spectroscopy required to measure [Ar III]7138 and the [S III] auroral line must be matched with broad wavelength coverage to detect the nebular [S III] lines and derive a reliable Te[S III].
@@
These abundance trends reveal that the typical galaxy in the CECILIA sample has lower S/O and Ar/O than the solar abundance ratio. We measure an average log(S/O) 1.780.21 dex and log(Ar/O) 2.640.13 dex (plotted as dot dashed red lines in Figure 8), which are 0.21 and 0.33 dex lower than the respective solar ratios. Consistent with local chemical abundance surveys, there is no apparent evolution in S/O or Ar/O with total metallicity. From CECILIA, there is growing evidence for non-solar abundance patterns in galaxies during the epoch of Cosmic Noon. Given the contribution of Type Ia SNe to the gas-phase S and Ar abundances, a natural explanation for the enhanced O relative to these elements is the preferential enrichment from CCSNe. We provide the S/O and Ar/O abundance ratios predicted from the IMF-averaged CCSNe yields for metal-poor (5% ) stars from kobayashi2006 as black dotted lines in Figure 8, which are log(S/O) 1.74 dex and log(Ar/O) 2.62 dex. These values are consistent with the minimum Ar/O predicted from chemical evolution models for the Milky Way (e.g., kobayashi2020) and M31 (arnaboldi2022; kobayashi2023). We observe that the average S/O and Ar/O measured in CECILIA is in good agreement with the yield predictions and minimum ratios expected from chemical evolution models, further supporting a scenario of significant CCSNe enrichment with minimal contributions from Type Ia SNe.
@@
The average ratios measured in the full sample are in good agreement with the abundance trends in D40, the first galaxy with direct Te[S III], S and Ar abundance measurements at z2 (rogers2024). As discussed in that work, certain systematic uncertainties could produce an offset to low S/O and Ar/O. First, D40 was previously missing a direct measurement of the low-ionization zone Te, such that the earlier O+ and S+ abundances could have been biased. From the combined MOSFIRE and NIRSpec data reported here, the Te[O II] in D40 agrees with the Te inferred from the other direct temperatures. While the lack of Te information did not appear to significantly bias the abundance ratios in D40, galaxies with a single Te are reliant on the functional form of the adopted Te scaling relations. Second, the selected ICFs could fail to account for all unobserved ionization states of S and Ar in the ISM. However, the trend of sub-solar S/O and Ar/O persists when the analysis is repeated with the amayo2021 ICFs (average S/O and Ar/O are 1.830.15 dex and 2.650.12 dex, respectively), indicating that a substantial change to the shape of the ICFs is required to bring the ratios into agreement with the solar abundances (see also stanton2025).
The use of the [O II] nebular lines from the MOSFIRE spectra could bias to larger O+ abundance (see discussion in previous subsection). However, when we repeat the analysis with only Te[S III] and use the [O II] auroral lines for the O+ abundance (similar to the original analysis for D40), we find that the S/O abundances in five of the six galaxies remain consistent within uncertainties. For BX628, using Te[S III] to infer the low-ionization zone Te produces log(S/O) 1.680.15 dex; this is in good agreement with the average S/O in CECILIA and may indicate that Te[O II] is biased high by the low-S/N MOSFIRE [O II] lines in this galaxy. Finally, we note that accounting for dust depletion would increase O in the most metal-rich galaxies (see §3.3). While the depletion of S onto dust grains is uncertain (jenkins2009), Ar is not depleted onto grains and the Ar/O abundance ratios in some of the CECILIA galaxies may be interpreted as a maximum if dust depletion is significant. In summary, while we cannot rule out the potentially distinct S or Ar ICFs, the average CECILIA S/O and Ar/O trends should be representative of the chemical abundance patterns in typical SFGs at Cosmic Noon.
In principle, S should be more accessible than Ar owing to the relatively bright [S II] and [S III] CELs in the optical and NIR: on average, the fluxes of [S III]9533 and [S II]6718 are a factor of 6.8 and 3.1 times larger than the flux of [Ar III]7138 in the same CECILIA galaxies. In practice, there are relatively few high-z galaxies with direct S/O abundances owing to the observational challenges associated with detecting the NIR [S III] lines. Instead, we can compare these abundance trends to local SFGs analogous to high-z galaxies in their SFRs. Recently, arellano-cordova2024a reported an average log(S/O) 1.730.10 dex in the highly star-forming galaxies of the CLASSY survey (berg2022; james2022). This is consistent with the average reported in the CECILIA galaxies (log(S/O)1.780.21 dex, or (S/O)0.620.30(S/O)⊙) and further supports a scenario where the gas-phase S abundance is sensitive to enrichment from CC and Type Ia SNe (see also esteban2025). However, the scatter in S/O observed in the CLASSY and CECILIA galaxies is large, and additional S abundances at high-z are necessary to assess the evolution of S/O over cosmic time.
@@
To compare with existing abundance measurements in Figure 9, we focus on Ar/O measured in other high-z galaxy surveys. We include Ar/O from the observations of SGAS 1723+34 from the TEMPLATES survey (welch2024), the Sunburst Arc (welch2025), MACSJ1149-WR1 (morishita2025), the EXCELS survey from stanton2025, and the compilation of Ar/O from bhattacharya2025A excluding D40 and SGAS 1723+34. We use the published Ar/O in each source except for MACSJ1149-WR1, which is reported to have log(Ar/O) 2.400.02 dex. The reported uncertainty on Ar/O is likely underestimated, considering that the uncertainty in the total O abundance is 0.1 dex and the low-ionization zone temperature has uncertainty in excess of 3000 K. We recompute the Ar/O abundance in MACSJ1149-WR1 using the reported line intensities and ionization zone temperatures in morishita2025 and find log(Ar/O) 2.390.22 dex, which we use when comparing to the CECILIA sample.
@@
Figure 9 shows log(Ar/O) vs. redshift for 36 galaxies at 1.2z5.8, where the literature abundances are plotted as purple triangles. The Ar/O abundances in the literature galaxies are generally at or below the solar abundance ratio. To assess the average trends of these high-z galaxies, we measure the average Ar/O ratio in three redshift bins: 1.5z2.5, 2.5z3.5, and 3.5z4.5. Each bin has more than five galaxies with direct abundances from which we measure the average Ar/O, which are plotted as white stars in Figure 9. While the highest redshift bin has the largest uncertainty, the galaxies at 1.5z2.5 and 2.5z3.5 show similar log(Ar/O) and are in agreement with the CCSNe yield predictions. Excluding the CECILIA sample, the number of galaxies in the 1.5z2.5 and 2.5z3.5 bins decrease from 15 to 5 and 12 to 4, respectively. Without the CECILIA galaxies, the average log(Ar/O) in these bins increases to log(Ar/O) 2.48 dex, consistent with the binned averages in Figure 9 and over 0.15 dex below the solar Ar/O ratio.